indicate the function of f(x) below in fourier series

"0,-1<x<0"

"f(x)=" {

"1,0<x<3"


1
Expert's answer
2021-05-26T02:20:06-0400

"T=4"


"a_0=\dfrac{1}{4}\displaystyle\int_{-1}^{3}f(x)dx=\dfrac{1}{4}\displaystyle\int_{0}^{3}dx"

"=\dfrac{1}{4}[x]\begin{matrix}
 3 \\
 0
\end{matrix}=\dfrac{3}{4}"

"a_n=\dfrac{1}{4}\displaystyle\int_{-1}^{3}f(x)\cos(\dfrac{n\pi x}{2})dt=\dfrac{1}{4}\displaystyle\int_{0}^{3}\cos(\dfrac{n\pi x}{2})dx"

"=\dfrac{1}{2n\pi}\big[\sin(\dfrac{n\pi x}{2})\big]\begin{matrix}
 3 \\
 0
\end{matrix}=\dfrac{1}{2n\pi}\sin\big(\dfrac{3n\pi }{2}\big)"



"n=1, a_{1}=-\dfrac{1}{2(1)\pi}"

"n=2, a_{2}=0"

"n=3, a_{3}=\dfrac{1}{2(3)\pi}"

"..."


"a_{2k+1}=\dfrac{(-1)^{k+1}}{2(2k+1)\pi}, k=0,1,2,..."


"b_n=\dfrac{1}{4}\displaystyle\int_{-1}^{3}f(x)\sin(\dfrac{n\pi x}{2})dx=\dfrac{1}{4}\displaystyle\int_{0}^{3}\sin(\dfrac{n\pi x}{2})dx"

"=-\dfrac{1}{2n\pi}\big[\cos(\dfrac{n\pi x}{2})\big]\begin{matrix}
 3 \\
 0
\end{matrix}=\dfrac{1}{2n\pi}\big(1-\cos\big(\dfrac{3n\pi }{2}\big)\big)"

"n=1, b_{1}=\dfrac{1}{2(1)\pi}"

"n=2, b_{2}=\dfrac{2}{2(2)\pi}"

"n=3, b_{3}=\dfrac{1}{2(3)\pi}"

"n=4, b_{4}=0"

"n=5,b_{5}=\dfrac{1}{2(5)\pi}"

"b_{6}=\dfrac{2}{2(6)\pi}"

"..."


"b_{2k+1}=\dfrac{1}{2(2k+1)\pi}, k=0,1,2,..."

"b_{4m+2}=\dfrac{2}{2(4m+2)\pi}, m=0,1,2,.."


"f(x)=\dfrac{3}{8}+\displaystyle\sum_{k=0}^{\infin}\dfrac{(-1)^{k+1}}{2(2k+1)\pi}\cos(\dfrac{(2k+1)\pi x}{2})"

"+\displaystyle\sum_{k=0}^{\infin}\dfrac{1}{2(2k+1)\pi}\sin(\dfrac{(2k+1)\pi x}{2})"




"+\displaystyle\sum_{m=0}^{\infin}\dfrac{1}{2(2m+1)\pi}\sin((2m+1)\pi x)"